The monthly revenue R achieved by selling x wristwatches is figured to be R(x)=71x-0.2x^2. The monthly cost C of selling wristwatches is C(x)=26x+1600.
a) How many wristwatches must the firm sell to maximize​ revenue? What is the maximum​ revenue?
b. Profit is given as P(x)=R(x)-C(x). What is a profit function?

Answer by Boreal(15235)

Maximum revenue maximizes -0.2x^2+71x, where a=-0.2 and b=71
the x-value for this quadratic's maximum if -b/2a=-71/-0.4 or 177.5 watches
177 revenue is
$18832.80
178 revenue is $6301.20
177 revenue is $6301.20
either of those maximizes revenue
profit function is their difference or -0.2x^2+45x-1600=P(x)